# Code-Driven Law NO, Normware SI!

### v1

@@ 4.4 Primary and secondary drivers
It is not in our scope to participate in the debate on what originates legal systems (eg. Kelsen’s Grundnorm, or Schmitt’s constituent power). Being in the computational realm provides us with a more controlled grounding than the real world, because some pre-existing regulatory infrastructure is necessarily given: at some micro-level, components will eventually behave as control systems, with no expectation of failure. These elementary capacities of artificial devices (physical, symbolic) are generally associated to powers granted to users/developers. From this insight, we can formulate a distinction between:
  • code-driven computational entities, whose control flow perfectly aligns with the instructions provided by another component or by a human with adequate powers (eg. the developer);
@@
Correspondingly, in terms of normware components, one can distinguish between primary drivers (eg. the ones hard-coded in the processor); from secondary drivers, which contingently contribute to further specify and constrain the system behaviour, in support to primary drivers. [^12] In principle, within secondary drivers, one could also envision a measure of normative strength or normative force, proportional to the level of influence of the code to the system’s behaviour. Deciding where the boundary lies between primary and secondary drivers is a fundamental design choice, which is obscured by traditional stances on computational artefacts.
@@ 5 Normware in the wild
When directives are implemented as primary drivers we are in the domain of code-driven and data-driven law. The concept of normware gains however most of its value once secondary drivers are introduced in the system. Differently from hardware and software, which are primarily defined in terms of control, we expect a piece of normware not to be used for control, but primarily for guidance of the computational system. This is a view that we apply naturally when we think of directives expressed to human agents (eg. parents to kids, legislators/policy-makers to social participants), but not to artificial devices. Yet, because normware abstracts the underlying computational technology, it can be used as a framework to evaluate any concept and architecture associated to computational systems: Turing machines, smart contracts, operating systems, the Internet, knowledge-based systems, neural networks and other machine learning methods (eg. ensemble methods), as well as more complex architectures. We will now review a few of these examples.
Figure 1: Examples of first-order control: eg. a user or developer provides a goal in the form of an instruction and the computer performs an action executing the associated operation. This association is hard-coded or logically derived in code-driven systems, or selected by some machine-learning method in data-driven systems.
@@ 5.1 First-order control
Algorithms expressed in terms of control flow of instructions (specifying the sequence of execution) can be reformulated in terms of control structure of computational actors [40]. [^13] But not always vice-versa, see eg. [40]. Consider for instance a “printing agent” implemented as a Turing machine. [^14] A Turing machine is an essential model of computation consisting of: a tape; a head, which can read/write symbols on the tape, and move the tape left/right; a state register; an action-table. Seen as an artificial agent, the head of a Turing machine maps to the agent’s physical body (sensors/actuators for reading/writing); the tape corresponds to the environment subjected to its physical actions; the state register corresponds to the mental state, whereas the action-table maps to automated behavioural responses. Suppose this agent is designed to print on its tape, sequentially and indefinitely, the symbol it observes in its state register, overwriting what is recorded on the tape. Now imagine that, assuming there is no mechanical impediment, a second agent (possibly human) reads/write the state register of the “printing agent”. The first machine, the controlled entity, can be seen as being subjected to directives (about printing) issued by the second one, the controller entity. The same pattern can be observed between developers and computational systems in code-driven and data-driven systems (see Fig. 1): these systems cannot escape from what they have been instructed to do by the developers. Control structures like this one have a one-to-one mapping to power relationships, and therefore can be captured by adequate normware specifications, although it may be add impractical overhead with simple hierarchical constructs.
@@ 5.2 Second-order control
Let us consider that the controller entity of the previous configuration is also a Turing machine. Its internal register may be modified in turn by a programmer, or by some other artificial entity enabled to so, and so on, realizing a hierarchy of control. The controller/controlled construct is the prototypical domain of study of (first-order) cybernetics. An additional control level brings us to the domain of second-order cybernetics. Similar three-levels views of systems are used in various disciplines, as for instance: operations, design/development, policy-making activities in organizations [41], or operational, tactical, strategic dimensions in decision-making (see Fig. 2).
In previous work [42], we argued that contemporary challenges of explainable AI can be framed as achieving reasonable abilities in justifying a certain decision, including being able to reject (ie. avoid to produce) unacceptable, “wrong” arguments. Similarly, trustworthy AI can be associated to the requirement of not taking “wrong” decisions, of performing “wrong” actions. In both cases, determination of unacceptability arises from conflicts between tactical and strategic levels, or between tactical and operational level. For instance, let us take the case of a computational model used for a classification task. Labels given with the training set are generally deemed to be adequately consistent, but in reality the rationale behind the labels may change in time. For instance, in administrative activities, the policies regulating the provided services may have changed, bringing to different categorization of similar cases. New behavioural trends may also have emerged in the population, which were not captured in the original training dataset, thus causing inaccuracies in prediction. Most common machine-learning architectures lack a proper diagnostics feedback for these type of misalignments: their architecture is essentially one of first-order control. However, even from theory, it is known that second-order feedback mechanisms are necessary for satisfying requirements of organizational intelligence (eg. Beer’s viable system model [43]). This suggests that reappraising the control structure of machine learning components within a larger normware network can be a valuable exercise (not too dissimilar to the “follow the money” principle). Indeed, answering to the question what is controlling what can help in identifying and then addressing deficiencies (for further examples see [42]).
@@ 5.2.1 Decentralizing control
Traditional programs, as well as inferential models constructed via machine learning, are generally conceptualized to be executed as single, unitary computational processes. Similarly, the knowledge level presented by Newell [9] assumes a single agent whose “law of behaviour” follows a “principle of rationality”. Or still, cybernetics: introduced as the study of regulation mechanisms “in the animal and the machine” (Wiener, [44]), even when approached in second-order forms, and applied to interpret collective entities as organizations, it maintains a fundamentally organicist view of systems. This stance has obvious limitations: ecological systems cannot generally be approached as organic entities.
In contrast, networking is plausibly the computational domain that realizes at best Deleuze and Guattari’s assemblage view of systems [6]. Even if most of the attention on “responsibility” goes nowadays to machine-learning-based AI, networking is also relevant to the responsible computing challenge, because, for the consequences it may cause or facilitate at individual and collective levels, transporting data is not a neutral endeavour. This is made clear for instance in the Responsible Internet proposal [45]. As argued in a subsequent elaboration [5], the distributed and decentralized nature of networks intrinsically requires instruments for:
  • dealing with a plurality of policies (of higher and lower abstractions, spatially and temporally located) set by users, network operators, or various governance bodies;
@@
The resulting view is a call for technologies enabling plural second-order control (see Fig. 3), whose participatory structures may resemble or put in continuity with those operating in human social systems.
Interestingly, the normware level elaborated in the previous sections provides a conceptual framework integrating these requirements in their entirety. The first two requirements concerns artifacts that, in normware terms, are meant to regulate behaviours, qualifications, and expectations, addressed in section 3. The third requirement is about institutional powers assigned to parties, encoding the backbone control structure of the infrastructure. The fourth requirement consists of meta-directives specifying mechanisms of conflict resolution, including anticipation. These last two were addressed in section 4, and captures what is needed to develop the technological counterpart of governance processes.
@@ 6 Research perspectives
Having set the conceptual horizon of normware, I will now briefly elaborate on two main research tracks relevant for “normware-centred” engineering: how to specify normware (for individual artefacts), and how to operationalize normware (as an ecological process), focusing in particular on conflicts.
@@ 6.1 Specifying code
Although the paper argues for going beyond a code-driven law perspective, code still plays a central role in code-guided law. To enable humans to specify normware in an intelligible and traceable way, we have plausibly to strive for some adequate mapping from (contextual interpretations of) sources of norms to some computational formal language. Since the first pioneering experiments with TAXMAN for US corporate tax law [12], and the modelling of the British Citizenship Act with a logic program [13], several solutions for specifying normative artifacts (norms, contracts, policies) in a computational processable way have been presented in the literature. Amongst the most recent efforts we acknowledge e.g. LegalRuleML [46, 47], PROLEG [48], InstAL [49], ODRL [50, 51], Symboleo [52], FLINT/eFLINT [53, 54], Logical English [55], Catala [56], Blawx [57], Stipula [58], and DPCL [59]. Legal core ontologies (eg. LKIF-core [60], UFO-L [61]) have also been proposed to systematize concepts and relationships relevant to normative reasoning. No solution amongst those has achieved general acceptance, even less in deployment. More unexpectedly, there exists no common ground (nor representational, nor computational) enabling a comparison between these solutions. Complementary to these efforts, there exist a number of industrial, domain-specific standards, which—in contrast to the above—are widely used, eg. XACML (for access control) [62], BGP policies (routing) [63], Protune, Rei, Ponder, TrustX (for cloud services) [64]. These solutions considers only few or even no explicit normative concepts; their concerns is not modelling what is in the law or in executive policies, but of having effective and computationally efficient instruments of regulation (for security, privacy, etc.). Similar considerations apply to artifacts to specify terminology and expectations like computational ontologies, logic programs, or functionally similar entities.
@@
Law-, policy-, and more in general knowledge-centred software is evaluated against both a language dimension (determining what types of knowledge we can specify and how), and a computational dimension (determining what we can actually do with these knowledge artifacts). Approaches giving priority to logical aspects result in well-defined semantics, which enables the possibility of using formal verification to guarantee the correctedness of the system. However, despite efforts in such sense, there are strong reasons to believe we may never reach an agreed “standard” semantics. First, there exist an endless number of semantic dimensions that play a role in law (deontic, potestative, action/dynamics, priority/defaults, intentionality). Standards seem a feasible objective only in narrow sub-domains (eg. primarily declarative, as for instance reasoning on the amount of due taxes), without capturing meta-norms and then overlooking at known properties of human norms like defeasibility. Not surprisingly, for reasons as reliability, reusability, and better performance, industrial formalisms have much simpler representational models. In many cases, one needs not to introduce explicit normative categories [65, 66, 26]. Second, any semantics chosen can be seen as a commitment, and any commitment can be argued and contested under certain conditions. Only hardware manufacturers are exempt from these considerations, as they have hard-coded physically (when correctly functioning) the local semantics of their components.
@@
Regardless of the formal verification aspect, this paper provided arguments for concluding that a central (and overlooked) problem in making norms executable is their social operationalization: ie. of transforming their declarative aspects into procedures, not in programming sense, but in the socio-legal sense: addressing governance mechanisms entails designing the right intervention points (who has the power to do what, and under which conditions).This aspect is also a matter of commitment. For instance, a duty can be operationalized with different types of enforcement, placing (or not placing) various burdens and incentives upon the social participants.
@@
Accepting that law-encoding software relies on choices of commitments—concerning both semantics and social operationalization—suggests also that these commitments should be (primarily) ascribed to the organizations using this software. Indeed, the presence of a strong standard (eg. issued by a central authority) may be used by parties to deny responsibility, and propagate systematic errors from central nodes. In these conditions, the only viable collective goal is to strive for a core integration standard, ie. a sort of knowledge interchange format, an agreement on a minimal informational model encompassing architectural aspects as the functions described in section 3 and 4.
@@ 6.2 Resolving conflicts
Dealing with pluralism means not only enabling the existence of several concurrent, potentially competing components, but also to have instruments to resolve such conflicts. Because these instruments are also coordination artifacts, they belong to the normware level. Indeed, one could encode in a normware artefact eg. the formal meta-rules determining the priority between laws, or voting mechanisms, or other relevant institutional patterns.
Relevant conflict resolution mechanisms have been researched in various disciplines, providing in principle off-the-shelf methods that can be reused in a normware-based computational system. For instance, aggregation of preferential structures, eg. by voting mechanisms, have been studied extensively in computational social choice (see eg. [67]), as well as other branches of applied mathematics [68]. Directives can also be seen as public arguments. Formal argumentation theory (eg. [69]) focuses on attacks (and/or support) relationships between individual arguments, and provides several methods to evaluate the strength of arguments based on the whole argumentation system. The various forms of defeasible logics introduce rule priorities in the derivation process, approaching conflict resolution in hierarchical terms; this is particularly relevant in the legislative context, for instance to model the hierarchy of law, as well to capture priorities between values underpinning norms. A completely different family of solutions for conflict resolution would consist of evolutionary computation techniques: a normware ecology may be seen as a population subjected to artificial selection, with fitness defined dually to conflict.
Whether or which of these techniques will be actually useful in practical settings is yet an open question, and out of our scope. Here it is relevant to observe that normware provides an unexpected ground of integration of efforts scattered across various computational sub-fields.
@@ 7 Conclusion
The concept of normware brings to the foreground that both institutional systems (particularly in their procedural aspects) and computational systems are in fact information processing systems. Indeed, any institutional system transcends its individual social participants, and, to the extent that such a system is designed (ie. its form has been deliberately chosen by rational agents), it can also be seen as an artificial one. Now, bureaucratic alienation in formal organizations (eg. public administrations) is a well known phenomenon, both for participants in these organizations (cf. Weber’s theory of bureaucratic management, business-IT alignment literature, etc.) and for the various stakeholders interacting with them. Computational systems can be related to an analogous, plausibly even more intense alienation.
Computers process symbols or numbers without any clue on their meaning further on in the decision process. This implies that, from a systematic point of view, no “intelligent” system can be expected without: (i) designing adequate feedback loops (and the associated intervention points); (ii) taking adequately into account a plurality of drivers and of contexts; (iii) being able to identify and overcome the conflicts that will naturally occur. These three points summarize the most distinctive matters we have elaborated in this paper, investigating the various aspects of normware. Engineering perspectives at an hardware (physical artefacts/processes), or software (symbolic artefacts/processes) levels do not provide the right abstraction to discuss about governance (coordination artefacts/processes) within the computational realm.
The continuity between institutional and computational systems however does not imply that accepting more sound computational forms of law will remove the need for human law. On the contrary, as human intervention is required to repair issues due to bureaucratic alienation, the best way to define their relationship (in the human and in the computational domain) is as the relationship higher-courts have towards lower-courts, ensuring that humans will always remain in control and mechanisms like appeals (against decisions) and quashing/overruling (against decision-making processes) become systemic. In fact, a core conclusion of the present contribution is that the socio-technical challenge we are facing is not mechanizing law, but introducing legitimate normative processes within the computational realm. Discussions on legitimacy do not solve challenges on technical possibility; normware may offer a common ground to discuss and advance the latter.

### v2

@@ 4.4 Primary and secondary drivers
It is not in our scope to participate in the debate on what originates legal systems (e.g. Kelsen’s Grundnorm, or Schmitt’s constituent power, see e.g. [40]). Being in the computational realm provides us with a more controlled grounding than the real world, because some pre-existing regulatory infrastructure is necessarily given: at some micro-level, components will eventually behave as control systems, with no expectation of failure. These elementary capacities of artificial devices (physical, symbolic) are generally associated to powers granted to users/developers. From this insight, we can formulate a distinction between:
  • code-driven computational entities, whose control flow perfectly aligns with the instructions provided by another component or by a human with adequate powers (e.g. the developer);
@@
Correspondingly, in terms of normware components, one can distinguish between primary drivers (e.g. the ones hard-coded by the manufacturer); from secondary drivers, which contingently contribute to further specify and constrain the system behaviour, on top of primary drivers. [^12] In principle, within secondary drivers, one could also envision some measure of normative strength or normative force, proportional to the level of influence of the directive to the actual system’s behaviour. Deciding where the boundary lies between primary and secondary drivers is a fundamental design choice, which is obscured by traditional stances on computational artefacts.
@@ 5 Normware in the wild
When directives are implemented as primary drivers we are in the domain of code-driven and data-driven law. The concept of normware gains however most of its value once secondary drivers are introduced in the system. Differently from hardware and software, which are primarily defined in terms of control, we expect a piece of normware not to be used for control, but primarily for guidance of the computational system. This is a view that we apply naturally when we think of directives expressed to human agents (e.g. parents to kids, legislators/policy-makers to social participants), but not to artificial devices.
Yet, because normware abstracts the underlying computational technology, it can be used as a framework to evaluate any concept and architecture associated to computational systems: Turing machines, smart contracts, operating systems, the Internet, knowledge-based systems, neural networks and other machine learning methods (e.g. ensemble methods), as well as more complex architectures. We will now review a few of these examples.
Figure 1: Examples of first-order control: e.g. a user or developer provides a goal in the form of an instruction and the computer performs an action executing the associated operation. This association is hard-coded or logically derived in code-driven systems, or selected by some machine-learning method in data-driven systems.
@@ 5.1 First-order control
From theory, it is know that algorithms expressed in terms of control flow of instructions—specifying the sequence of execution—can be reformulated in terms of control structure of computational actors (but not vice-versa) [41]. Consider for instance a “printing agent” implemented as a Turing machine. [^13] A Turing machine is an essential model of computation consisting of: a tape; a head, which can read/write symbols on the tape, and move the tape left/right; a state register; an action-table. Seen as an artificial agent, the head of a Turing machine maps to the agent’s physical body (sensors/actuators for reading/writing); the tape corresponds to the environment subjected to its physical actions; the state register corresponds to the mental state, whereas the action-table maps to automated behavioural responses. Suppose this agent is designed to print on its tape, sequentially and indefinitely, the symbol it observes in its state register, overwriting what is recorded on the tape. Now imagine that, assuming there is no mechanical impediment, a second agent (possibly human) reads/write the state register of the “printing agent”. The first machine, the controlled entity, can be seen as being subjected to directives (about printing) issued by the second one, the controller entity. The same pattern can be observed between developers and computational systems in code-driven and data-driven systems (see Fig. 1): these systems cannot escape from what they have been instructed to do by the developers. Note how control structures like this one have a one-to-one mapping to power relationships, and therefore can be captured by adequate normware specifications, although this effort adds impractical overhead with simple hierarchical constructs as the one expressed in this example.
@@ 5.2 Second-order control
Let us consider that the controller entity of the previous configuration is also a Turing machine. Its internal register may be modified in turn by a programmer, or by some other artificial entity enabled to so, and so on, realizing a hierarchy of control. While the controller/controlled construct is the prototypical domain of study of (first-order) cybernetics, an additional control level brings us to the domain of second-order cybernetics. Similar three-levels views of systems are however introduced in other domains, as for instance: operations, design/development, policy-making activities in organizations [42], or operational, tactical, strategic dimensions in decision-making (see Fig. 2).
In previous work [43], we argued that contemporary challenges of explainable AI can be framed as achieving reasonable abilities in justifying a certain decision, including being able to reject unacceptable, “wrong” arguments. Similarly, trustworthy AI can be associated to the requirement of not taking “wrong” decisions, of performing “wrong” actions. In both cases, determination of unacceptability arises from conflicts between tactical and strategic levels, or between tactical and operational level. As an example, let us take the case of a computational model used for a classification task. Labels given with the training set are generally deemed to be consistent, but in reality this may be not the case. For instance, in administrative activities, the policies regulating the provided services may change, bringing to different categorization of similar cases. New behavioural trends may also emerge in the population, which were not captured in the original training dataset, thus causing inaccuracies in prediction. Most common machine-learning architectures lack a proper diagnostics feedback for these type of misalignments: their architecture is essentially one of first-order control. However, it is known that second-order feedback mechanisms are necessary for satisfying requirements of organizational intelligence (e.g. Beer’s viable system model [44]). This suggests that reappraising the control structure of machine learning components within a larger normware network can be a valuable exercise. Answering to the question what is controlling what can help in identifying and then addressing deficiencies (for further examples see [43]).
@@ 5.2.1 Decentralizing control
Traditional programs, as well as inferential models constructed via machine learning, are generally conceptualized to be executed as single, unitary computational processes. Similarly, the knowledge level presented by Newell assumes a single agent whose “law of behaviour” follows a “principle of rationality” [9]. Or still, cybernetics. Introduced as the study of regulation mechanisms “in the animal and the machine” (Wiener, [45]), even when approached in second-order forms, and applied to interpret collective entities as organizations, cybernetics maintains a fundamentally organicist view of systems. This stance has obvious limitations, as it cannot support adequately the understanding of ecological systems (in our case, complex computational architectures).
In contrast, networking is plausibly the computational domain that realizes at best Deleuze and Guattari’s assemblage view of systems [6]. Even if most of the attention on “responsible computing” goes nowadays to machine-learning-based AI (and in particular generative AI, including chatbots based on LLMs), networking is also relevant to this challenge, because transporting data is not a neutral endeavour, for the consequences it may cause or facilitate at individual and collective levels. This is made clear for instance in the Responsible Internet proposal [46]. As argued in a subsequent elaboration [5], the distributed and decentralized nature of networks intrinsically requires instruments for:
  • dealing with a plurality of policies (of higher and lower abstractions, spatially and temporally located) set by users, network operators, and/or the various governance bodies;
@@
The resulting view is a call for technologies enabling plural second-order control (see Fig. 3), whose participatory structures may resemble, if not put in continuity, with those operating in human social systems.
Interestingly, the normware level provides a conceptual framework integrating these requirements in their entirety. The first two requirements concerns artifacts that are meant to regulate behaviours, qualifications, and expectations, addressed in section 3. The third requirement is about institutional powers assigned to parties, encoding the backbone control structure of the infrastructure. The fourth requirement consists of meta-directives specifying mechanisms of conflict resolution, including anticipation. These last two were addressed in section 4, and capture what is needed to develop a computational counterpart of governance processes.
@@ 6 Research perspectives
Having set the conceptual horizon of normware, we will now briefly elaborate on two main research tracks relevant for a “normware-centred” engineering: how to specify normware (for individual artefacts), and how to operationalize normware (as an ecological process), focusing in particular on conflicts.
@@ 6.1 Specifying code
Although the paper argues for going beyond a code-driven law perspective, code still plays a central role in code-guided law. To enable humans to specify normware in an intelligible and traceable way, we have plausibly to strive for some adequate mapping from (contextual interpretations of) sources of norms to some computational formal language. Since the first pioneering experiments with TAXMAN for US corporate tax law [12], and the modelling of the British Citizenship Act with a logic program [13], several solutions for specifying normative artifacts (norms, contracts, policies) in a computational processable way have been presented in the literature. Amongst the most recent efforts we acknowledge e.g. LegalRuleML [47, 48], PROLEG [49], InstAL [50], ODRL [51, 52], Symboleo [53], FLINT/eFLINT [54, 55], Logical English [56], Catala [57], Blawx [58], Stipula [59], and DPCL [60]. Legal core ontologies (e.g. LKIF-core [61], UFO-L [62]) have also been proposed to systematize concepts and relationships relevant to normative reasoning. No solution amongst those has achieved general acceptance, even less in deployment. More unexpectedly, there exists no common ground (nor representational, nor computational) enabling a comparison between these solutions. Complementary to these efforts, there exist a number of industrial, domain-specific standards, which—in contrast to the above—are widely used, e.g. XACML (for access control) [63], BGP policies (for routing in networking) [64], Protune, Rei, Ponder, TrustX (for cloud services) [65]. These solutions considers only few or even no explicit normative concepts; their concerns is not modelling what is in the law or in executive policies, but of having effective and computationally efficient instruments of regulation (for security, privacy, etc.). Similar considerations apply to artifacts to specify terminology and expectations like computational ontologies, logic programs, or functionally similar entities.
@@
Law-, policy-, and more in general knowledge-centred software is evaluated against both a language dimension (determining what types of knowledge we can specify and how), and a computational dimension (determining what we can actually do with these knowledge artifacts). Approaches giving priority to logical aspects result in well-defined semantics, which enables the possibility of using formal verification to guarantee the correctedness of the system. However, despite efforts in such sense, there are strong reasons to believe we may never reach an agreed “standard” semantics. First, there exist an endless number of semantic dimensions that play a role in law (deontic, potestative, action/dynamics, priority/defaults, intentionality). Standards seem a feasible objective only in narrow sub-domains (e.g. primarily declarative, as for instance reasoning on the amount of due taxes), without capturing meta-norms and then overlooking at known properties of human norms like defeasibility. Not surprisingly, for reasons as reliability, reusability, and better performance, industrial formalisms have much simpler representational models. In many cases, one needs not to introduce explicit normative categories [66, 67, 26]. Second, any semantics chosen can be seen as a commitment, and any commitment can be argued and contested under certain conditions (as we already observed in section 2.3). Only hardware manufacturers are exempt from these considerations, as they have hard-coded physically the local semantics of their components.
@@
Besides the formal verification aspect, a central (and overlooked) problem in making norms executable is their social operationalization: i.e. of transforming their declarative aspects into procedures. Not in the programming sense, but in the socio-legal sense: addressing governance mechanisms entails designing the right intervention points (who has the power to do what, and under which conditions.This aspect is also a matter of commitment. For instance, a duty can be operationalized with different types of enforcement, each placing (or not placing) different burdens and incentives upon the social participants.
@@
Accepting that law-encoding software relies on choices of commitments—concerning both semantics and social operationalization—suggests also that these commitments should be (primarily) ascribed to the organizations using this software. The presence of a strong standard (e.g. issued by a central authority) may be used by parties to deny responsibility, and propagate systematic errors from central nodes. In these conditions, the only viable collective goal is to strive for a core integration standard, i.e. a sort of knowledge interchange format, an agreement on a minimal informational model encompassing architectural aspects as the functions described in section 3 and 4.
@@ 6.2 Resolving conflicts
Dealing with pluralism means not only enabling the existence of several concurrent, potentially competing components, but also to have instruments to resolve such conflicts. Because these instruments are also coordination artifacts, they belong to the normware level too. Indeed, one could encode in a normware artefact e.g. the formal meta-rules determining the priority between laws, or voting mechanisms, or other relevant institutional patterns.
Relevant conflict resolution mechanisms have been researched in various disciplines, providing in principle off-the-shelf methods that can be reused in a normware-based computational system. For instance, aggregation of preferential structures, e.g. by voting mechanisms, have been studied extensively in computational social choice (see e.g. [68]), as well as other branches of applied mathematics [69]. Directives can also be seen as public arguments. Formal argumentation theory (e.g. [70]) focuses on attacks (and/or support) relationships between individual arguments, and provides several methods to evaluate the strength of arguments based on the whole argumentation system. The various forms of defeasible logics introduce rule priorities in the derivation process, approaching conflict resolution in hierarchical terms; this is particularly relevant in the legislative context, for instance to model the hierarchy of law, as well to capture priorities between values underpinning norms. Defeasible reasoning accounts of precedent are also studied in the case-based reasoning research track [71, 72]. However, the ecological perspective of normware opens also to a completely different way to look at the problem, via evolutionary computation techniques: a normware ecology may be seen as a population subjected to artificial selection, with fitness defined dually to conflict.
Which of these techniques will be actually useful in practical settings (and under which conditions) is clearly an open question, and out of the scope of this paper. Yet, it is relevant to observe we have found an unexpected ground of integration for independent efforts scattered across various computational sub-fields.
@@ 7 Conclusion
The concept of normware brings to the foreground that both institutional systems (particularly in their procedural aspects) and computational systems are in fact information processing systems. Indeed, any institutional system transcends its individual social participants, and, to the extent that such a system is designed (i.e. its form has been deliberately chosen by rational agents), it can also be seen as an artificial one.
Now, bureaucratic alienation in formal organizations (e.g. public administrations) is a well known phenomenon, both for participants in these organizations and for the various stakeholders interacting with them. Computational systems can be related to an analogous, plausibly even more intense alienation. Computers process symbols or numbers without any clue on their meaning further on in the decision process.
From a systematic point of view, no “intelligent” system can be expected without: (i) designing adequate feedback loops (and the associated intervention points); (ii) taking adequately into account a plurality of drivers and of contexts; (iii) being able to identify and overcome the conflicts that will naturally occur. These three points summarize the most distinctive matters we have elaborated in this paper, investigating the various aspects of normware. Engineering perspectives at an hardware (physical artefacts/processes), or software (symbolic artefacts/processes) levels do not provide the right abstraction to discuss about governance (i.e. coordination artefacts/processes) within the computational realm.
The continuity between institutional and computational systems however does not imply that accepting more sound computational forms of law will remove the need for human law. On the contrary, as human intervention is required to repair issues due to bureaucratic alienation, the best way to define their relationship (in the human and in the computational domain) is as the relationship that higher-courts have towards lower-courts, ensuring that humans will always remain in control and mechanisms like appeals (against decisions) and quashing/overruling (against decision-making processes) become systemic.
In fact, a core conclusion of the present contribution is that the socio-technical challenge we are facing is not mechanizing law, but introducing legitimate normative processes within the computational realm. Discussions on legality and legitimacy cannot solve the challenges on designing/developing technical possibilities supporting both; normware-centred discussions are instead required to advance the latter dimension.